The width of a triangle is one-fifth as much as the length. The perimeter is 120cm. Find the length and the width.

Answer by fractalier(6550)

Perimeter P = 2L + 2W = 120 cm
W = (1/5)L or
L = 5W
Now substitute this in and get
2(5W) + 2W = 120
12W = 120
W = 10 cm
L = 50 cm